Billy invested $18 dollars of savings in two major business enterprises which paid interest at the rate of 3% and 6%, respectively. If his annual income for these two enterprises was equivalent to a return of 5% on the entire investment, find how much he invested at 3%.

Answer by Boreal(15235)

x at 3%
18-x=6%
.03x+.06(18-x)=18*0.05=0.9
.03x+1.08-.06x=0.9
-0.03x=-0.18
x=$6 *0.03=0.18 ANSWER
18-x=$12*0.06=0.72
they add to 0.90, which is 5% of $18.